OpenStudy (anonymous):

Write the point-slope form of the equation of the line through the given point with the given slope. through: (−3, 1), slope = − 1/2

OpenStudy (anonymous):

y-ysub1 = m ( x-xsub1) y-1=1/2(x-(-3) y-1=1/2x+3/2 y=1/2x+3/2+1 y=1/2x+5/2

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

That is the Formula For Point slope Form.